OpenStudy (anonymous):

Simplify e^(lnx^2-3lny). Does e cancel out only one ln or both in this problem?

OpenStudy (anonymous):

e^(lnx^2-3lny). = e^[ ln (x^2 / y^3) ] = x^2 / y^3

OpenStudy (anonymous):

Oh okay. I get it. Thank you!